Find the equation of the hyperbola with vertices (1,−4) and (1,4), and foci at
(1,−6) and (1,6)
1
Expert's answer
2020-03-02T11:27:55-0500

One thing to notice in all the coordinates is that their x coordinate are same i.e. all the points lie on a line x=1

hence the center of the hyperbola = (1,(-4+4)/2) = (1,0)

distance between center and vertex (a) = 4

the foci are at distance 6 from the center , hence c=6

"a^2+b^2=c^2
\newline b^2=20"

the equation of the hyperbola will be

"(y-k)^2/a^2 - (x-h)/b^2=1 \newline
(y-0)^2/16 -(x-1)/20 =1"